OpenStudy (loser66):

If p is prime, then \(n= 2^{p-1}(2^p -1)\) is perfect. True or False. Please, help

ganeshie8 (ganeshie8):

The statement is clearly false. Below is true however : If \(2^n-1\) is prime, then \(2^{n-1}(2^n-1)\) is a perfect number.

OpenStudy (loser66):

We should prove it, right?

ganeshie8 (ganeshie8):

which statement you wanto prove ?

OpenStudy (loser66):

Your statement.

ganeshie8 (ganeshie8):

simply evaluate the sum of divisors and see

ganeshie8 (ganeshie8):

Let \[m = 2^{n-1}(2^n-1)\] since \(2^n-1\) is prime, the numbers \(2^{n-1}\) and \(2^n-1\) are relatively prime, yes ?

OpenStudy (loser66):

yes

ganeshie8 (ganeshie8):

use the fact that \(\sigma \) is multiplicative

ganeshie8 (ganeshie8):

\(m = 2^{n-1}(2^n-1)\) \(\sigma(m) = \sigma(2^{n-1}(2^n-1)) = \sigma(2^{n-1})*\sigma(2^n-1)\)

OpenStudy (loser66):

hold on. I don't know what \(\sigma\) is. I saw it many times when searching practice problem online , but I was not taught what it is yet.

ganeshie8 (ganeshie8):

Ohk, we don't really need to know about \(\sigma \) here

ganeshie8 (ganeshie8):

but we need to find some way to find the sum of divisors of an integer

ganeshie8 (ganeshie8):

suppose \(m = 2^3*3\) can you find the sum of divisors of \(m\) ?

OpenStudy (pawanyadav):

Is it (1+2+3+4+6+8+12+24)

OpenStudy (dan815):